# RealMAN: A Real-Recorded and Annotated Microphone Array Dataset for Dynamic Speech Enhancement and Localization

### v1

@@ 5       Initialize a score map;
7             Within the score map, calculate the pixel score based on color intensity and brightness;
@@
10      Find the pixel of maximum score in the score map, as the detected position of LED light (source source);
@@ D.1 Details for experimental configuration
(Section 4.2) All speech enhancement and sound source localization networks are trained on 16 kHz 4-second speech clips, and tested on 16 kHz speech utterances.
For speech enhancement networks, their training configurations are set the same as in their original papers. Each model is trained for 50 epochs, and the best checkpoint is selected according to the DNSMOS score of validation data.
For the sound source localization task, the window length of STFT is 512 samples (32 ms) with a frame shift of 256 samples (16 ms). The model outputs a localization result every 6 frames. The Adam is used as the optimizer for training. The batch size of the fixed-array model and variable-array model are set to 16 and 4, respectively. The learning rate is initially set to 0.0005, and exponentially decays with a decaying factor of 0.975. We train each model for 50 epochs and the best checkpoint is selected according to the validation loss.
@@ D.2 Analysis of the spatial coherence of real multi-channel noise
One of the major differences between simulated diffuse noise and real-world recorded noise manifests as the spatial coherence [1]. We show the spatial correlation coefficient of the noise signals recorded with a 6cm-spaced microphone pair in some representative scenes, and the theoretical spherical (3D diffuse) noise field in Fig. 5. It can be seen that the spatial correlation of real noise largely varys from scene to scene. Some scenes, such as the ‘Market’, have a noise field that is fairly close to the 3D diffuse noise field. However, for many other scenes, the noise field is a complicated combination of directional noise sources and (partially) diffuse noise sources, and thus largely deviate from the 3D diffuse noise field. To demonstrate the temporal stationarity in terms of the spatial correlation of real noise, we plot the curve of spatial correlation as a function of time in Fig. 6, where the spatial correlations are estimated from consecutive 1-second recordings. This figure shows that, even for the ‘Market’ scene that has an overall 3D diffuse noise field, the spatial correlation is still highly time varying.
(a) Market.
(b) LivingRoom.
(c) Park.
(d) OfficeLobby.
Figure 5: Spatial correlation coefficients of multi-channel noise in typical scenes.
Figure 6: The (absolute value of) spatial correlation as a function of time, at the frequency of 2 kHz in the ’Market’ scene.
@@ D.3.1 Benchmark experiments
In addition to the low-SNR conditions presented in Section 4.2 where the recorded speech and noise are added, in this section, we present the results for high-SNR conditions where only the recorded speech (without adding extra noise) are tested. Table 10 and 11 show the performance of speech enhancement and sound source localization, respectively. Under the high-SNR conditions, noise is no longer a crucial factor, and accordingly the networks trained with real noise and simulated noise achieve comparable performance measures.
@@ D.3.2 Variable-array networks and array generalization
unprocessed - - 1.35 -4.7 2.58 2.36 1.99 10.1 1.21 -4.9 2.28 2.00 1.70 10.5
FaSNet-TAC [10] sim sim 1.74 -0.3 2.95 3.74 2.61 13.3 1.62 -0.4 2.86 3.63 2.49 14.2
sim real 1.89 -0.0 3.04 3.69 2.67 11.2 1.72 0.1 2.94 3.57 2.54 12.2
real sim 1.91 4.7 2.93 3.75 2.60 12.9 1.68 2.6 2.79 3.59 2.42 14.6
real real 1.97 4.6 2.99 3.75 2.65 12.0 1.73 2.5 2.87 3.62 2.50 12.9
SpatialNet [28] sim sim 1.97 -2.7 3.39 3.30 2.74 9.1 1.76 -2.7 3.35 3.16 2.65 9.8
sim real 1.84 -0.1 3.05 3.00 2.41 10.4 1.70 -0.1 2.98 2.84 2.30 11.2
real sim 2.74 7.8 3.24 3.66 2.82 8.9 2.39 5.3 3.15 3.53 2.68 9.8
real real 2.71 8.0 3.22 3.78 2.87 9.0 2.35 5.1 3.13 3.71 2.75 10.4
@@
sim sim 84.5 4.3 84.3 2.8
sim real 87.8 4.3 85.4 3.8
real sim 94.4 3.6 91.3 2.3
real real 92.7 3.5 91.4 2.3
@@
unprocessed 1.35 -4.7 2.58 2.36 1.99 10.1 1.21 -4.9 2.28 2.00 1.70 10.5
Fixed-Array 1.87 4.1 2.95 3.77 2.63 13.3 1.66 2.0 2.83 3.64 2.47 15.5
Variable-Array 1.73 3.0 2.84 3.68 2.50 15.5 1.58 1.6 2.73 3.54 2.35 18.1
@@
Fixed-Array[12] 86.8 2.9 93.8 1.7
Variable-Array[12] 86.3 3.2 85.7 2.7

### v2

@@ 5       Initialize a score map;
7             Within the score map, calculate the pixel score based on the color intensity and brightness;
@@
10      Find the pixel of maximum score in the score map as the detected position of the LED light (sound source);
@@ D.1 Details for experimental configuration
All speech enhancement and sound source localization networks are trained on 16 kHz 4-second speech clips, and tested on 16 kHz varying-long speech utterances.
For speech enhancement networks, the training configurations are set the same as in their original papers. Each model is trained for 50 epochs, and the best checkpoint is selected according to the DNSMOS score of the validation set.
For the sound source localization task, the window length of STFT is 512 samples (32 ms) with a frame shift of 320 samples (20 ms). The model outputs one localization result for every 5 frames. The Adam is used as the optimizer for training. The batch size of the fixed-array model and variable-array model are set to 16 and 4, respectively. The learning rate is initially set to 0.0005, and exponentially decays with a decaying factor of 0.975. We train each model for 50 epochs and the best checkpoint is selected according to the validation loss.
@@ D.2 Analysis of the spatial coherence of real multi-channel noise
One of the major differences between simulated diffuse noise and real-world recorded noise manifests as the spatial coherence [1]. We show the spatial correlation coefficient of the noise signals recorded with a 6cm-spaced microphone pair in different scenes, and the theoretical spherical (3D diffuse) noise field in Fig. 9. It can be seen that the spatial correlation of real noise largely varies from scene to scene. Most of the indoor and transportation scenes have a noise field that is fairly close to the 3D diffuse noise field, with several exceptions such as LivingRoom1 and OfficeLoby. In semi-outdoor and outdoor scenes, ambient noise is a complicated and dynamic combination of (partially) directional noise sources and (partially) diffuse noise sources from open space, and correspondingly their noise field generally deviate the 3D diffuse noise field.
To demonstrate the temporal stationarity in terms of the spatial correlation of real noise, in Fig. 10. we plot the curve of spatial correlation as a function of time for the ‘Market’ scene, where the spatial correlations are estimated from consecutive 1-second recordings. This figure shows that, even for the ‘Market’ scene that has an overall 3D diffuse noise field, the spatial correlation is still highly time varying and non-stationary.
(a) Indoor scenes (part1).
(b) Indoor scenes (part2).
(c) Indoor scenes (part3).
(d) Transportation scenes.
(e) Semi-outdoor scenes.
(f) Outdoor scenes.
Figure 9: Spatial correlation coefficients of multi-channel noise in different scenes.
Figure 10: The (absolute value of) spatial correlation as a function of time, at the frequency of 2 kHz for the ’Market’ scene.
@@ D.3.1 Benchmark experiments
In addition to the low-SNR conditions presented in Section 4.2 where the recorded speech and noise are added, in this section, we present the results for high-SNR conditions where only the recorded speech (without adding extra noise) are tested. Table 10 and 11 show the performance of speech enhancement and sound source localization, respectively.
@@ D.3.2 Variable-array networks and array generalization
unprocessed - - 1.34 -4.8 2.59 2.38 2.00 9.9 1.20 -4.9 2.28 2.00 1.70 10.3
FaSNet-TAC [13] sim sim 1.74 -0.3 2.96 3.63 2.58 12.7 1.62 -0.2 2.86 3.46 2.42 14.0
sim real 1.84 0.3 3.03 3.69 2.67 11.0 1.67 0.3 2.93 3.54 2.53 11.0
real sim 1.81 4.3 2.90 3.79 2.60 13.3 1.62 2.3 2.75 3.67 2.42 13.5
real real 1.83 4.3 2.97 3.72 2.63 12.0 1.64 2.2 2.87 3.57 2.49 11.7
SpatialNet [34] sim sim 2.03 -0.9 3.39 3.48 2.84 8.9 1.81 -0.8 3.37 3.34 2.75 10.0
sim real 1.22 -6.0 1.81 2.09 1.50 21.6 1.19 -5.8 1.69 1.90 1.40 26.2
real sim 2.69 8.0 3.28 3.68 2.87 8.9 2.33 5.2 3.20 3.55 2.73 9.9
real real 2.66 8.6 3.32 3.69 2.90 8.9 2.33 5.2 3.24 3.62 2.80 10.6
@@
sim sim 75.2 3.1 74.5 3.5
sim real 64.8 10.3 59.7 9.6
real sim 92.7 1.8 92.8 2.1
real real 90.9 1.8 92.8 2.0
@@
unprocessed 1.34 -4.8 2.59 2.38 2.00 9.9 1.20 -4.9 2.28 2.00 1.70 10.8
Fixed-Array 1.77 4.2 2.94 3.70 2.59 13.7 1.59 1.5 2.83 3.54 2.43 15.6
Variable-Array 1.75 3.5 2.92 3.75 2.59 14.1 1.58 1.3 2.78 3.59 2.42 16.1
@@
Fixed-Array[15] 89.0 2.1 91.4 2.1
Variable-Array[15] 88.3 2.3 84.3 2.9